. Hydrogen has a density of 89.9 grams per cubic meter at a certain temperature and a certain pressure. Calculate the volume in cubic meters of hydrogen having a mass of 0.1 kilogram at the same temperature and pressure.
1
Expert's answer
2019-12-16T11:06:44-0500

We write the Mendelev-Clapeyron equation for the first case

"PV_1=\frac{m_1}{\mu}RT"

we write the equation in another form

"P=\frac{m_1}{\mu V_1}RT"

Where

"\frac{m_1}{V_1}=\rho_1=89.9{g/m^3}=0.0899{kg/m^3}"

then we get the equation

"P=\frac{\rho_1}{\mu}RT" (1)

We write the Mendelev-Clapeyron equation for the second case

"PV_2=\frac{m_2}{\mu}RT" (2)

divide equation (2) by equation (1)

"\frac{PV_2}{P}=\frac{\frac{m_2}{\mu}RT}{\frac{\rho_1}{\mu}RT}"

we reduce such terms, we get

"\frac{V_2}{1}=\frac{m_2}{\rho_1}"

or "V_2=\frac{m_2}{\rho_1}=\frac{0,1}{0,0899}=1.112m^3"